i have a problem with a question asked on my homework. the instructions say; write the ratio of the first measurement to the second measurement. write both measurements in the same unit..... the first problem is; 84 ft, 54 in. i know about ratios but i dont understand this one. :(
Found 2 solutions by mananth, Alan3354:
Answer by mananth(16946)

Whenever you are taking a ratio both the quantities should have the same units
84ft = 84 * 12= 1008 in
54 in
1008/54 =168/9

Answer by Alan3354(69443)

i have a problem with a question asked on my homework. the instructions say; write the ratio of the first measurement to the second measurement. write both measurements in the same unit..... the first problem is; 84 ft, 54 in.
-----------------------
Change ft to inches
84ft = 84 * 12 = 1008 in
54 in
1008/54 = 56/3